APPLY THE CONCEPTS: Use the CVP graph to analyze the effects of changes in price and costs Graph the following on your own paper. At the original position, the break-even point in sales dollars is $24,000 at 500 units. The fixed costs are $8,000. Assume the slope of the sales line is equal to the selling price. When the two points of the sales line are at the origin and the break-even point, you see that the slope of the line is $48, which means that the selling price is $48. When the two points of the total costs line are at the origin and the break-even point, you see that the slope of the line is $32.00, which means that the variable cost per unit is $32. Leave the break-even point (x) at its original position. Use it as a reference point to answer the following questions. Analyze the scenarios by sliding the points on the lines to get the slope desired. Recall that the new break-even point for each scenario exists where the sales and total costs lines intersect. Compare it to the original break-even point (x). (You may want to put the lines back to their original position for each scenario.) Each scenario should be considered independently. 1. The company purchases a fixed asset and increases fixed costs by $2,000. Variable costs remain the same, which means that the slope does not change. This will cause the break-even point to move to the right, which means that break-even point in sales dollars increases. 2. A new supplier can provide a product of equal quality at $4.00 per unit less than the current direct materials cost. If the new supplier is used, the slope of the total costs line will be $, and the break-even point in sales dollars decreases. Can someone assist with #2, what is the slope of the total costs line?
Answer #1
Original break even point = Fixed Cost / Contribution margin
Contribution margin = (Sales - Variable)/ Sales = (48 - 32)/48 = 16/48 = 33.34%
Original break even point in sales dollars= 8000 / 33.34% = $23995.2
Req. 1) Increase in FC by $2000
New Fixed Cost = $8000 + $2000 = $10000
New break even point in sales dollars= 10000 / 33.34% = $29994
Req. 2) If new supplier provides RM at $4 less, then
Variable cost per unit would be $32 - $4 = $28
The slope of the total costs line will be $28.
Contribution margin = (48-28)/48 = 20/48 = 41.67%
Revised break even point in sales dollars= 8000 / 41.67% = $19198.46
The slope of the total costs line is the variable part of the cost exist in the total costs.
The fixed portion of the total costs is constant so total costs does not
differ due to that, so variable portion is the cost affecting part or its slope.